OpenStudy (kirbykirby):

How to prove "If a|15, then a|30" | is "divides"

OpenStudy (anonymous):

a|15 means there exists an x such that x*a = 15, and because 30 = 2*15, you get 2*x*a = 2*15=30, therefore, a|30

OpenStudy (anonymous):

human?

OpenStudy (anonymous):

:P Panda?

OpenStudy (anonymous):

"a divides 15" means you can multiply a by a certain number x to get 15, and as 30 is two times 15, if you multiply a by two times that x you get 30, see? So, there exists a number so that a times that number is 30, and that is the definition of "divides"

OpenStudy (kirbykirby):

thxx =]